find the length of a rectangle if it is four less than twice the width and the perimeter is 40 cm
Answer by Ahha(5)

Assign a variable to one of the sides. It's easier to assign it to the width because we can the do "four less than twice the width"
Let x be the width
Length(four less than twice the width) is 2x-4
The perimeter is what you get when you add up ALL of the exterior sides. In a rectangle there are four sides. We add up the expressions for all four sides and make it equal to 40. We don't need brackets when adding but it's good practice.
(x)+(2x-4)+(x)+(2x-4)=40
x+2x-4+x+2x-4=40 remove brackets
6x-8=40 Collect terms
6x=48 Add 8 to both sides
x=8 divide both sides by 6
This is the value of x which is the width.
To get the length you must plug it back in to the length expression.
2x-4
=2(8)-4
=16-4
=12
So the width is 12cm and the length is 8cm.
Check: Do these sides add up to the perimeter? 12+8+12+8=40. Yes they do...